Derive Schrodinger’s wave equation for a free particle in free space


1
Expert's answer
2021-03-04T11:54:23-0500

The operator for kinetic energy is:


"p = -i\hbar\dfrac{\partial}{\partial x}"

and for full energy is:


"E = i\hbar\dfrac{\partial}{\partial t}"

Using the analogy with the classical case (where for a free particle "E = \dfrac{p^2}{2m}") and applying these operators to a wavefunction "\Psi", obtain the Schrodinger’s wave equation for a free particle:


"-\dfrac{\hbar^2}{2m}\dfrac{\partial^2 \Psi}{\partial x^2} = i\hbar\dfrac{\partial\Psi}{\partial t}"



More detailed derivation see hee: http://hyperphysics.phy-astr.gsu.edu/hbase/quantum/Schr2.html.